hockey12151454:

how do you find the slope in y=3x+5 @summerdonovan149

hockey12151454:

help @summerdonovan149

Slumpedk:

why you want her help

Florisalreadytaken:

\[ y=mx+b \] from where \[ m=slope \]

Florisalreadytaken:

compare \[ y=mx+b \] \[ y=3x+5 \] m=3 that'd be your slope :D

carmelle:

The slope intercept form is \[y = mx +b\] where m = slope and b is y-intercept Based on this information, what do u think the slope is?